Aggregate Inc. plans to sell 10,000 heating units to a customer, and the cost per unit to Aggregate is $800. This cost is 20% variable and 80% fixed. Approximately how much will the cost per unit be to Aggregate if the customer decides to buy only 9,000 units?
Answer #1

Variable cost per unit=800*20%=$160

Fixed cost per unit=800*80%=$640

Total fixed costs =(640*10,000)=$6,400,000

Hence fixed cost per unit for 9000 units would be=(6,400,000/9000)=$711.11 per unit(Approx)

Hence cost per unit for 9000 units=Variable cost+Fixed cost

=(711.11+160)

=$871.11 per unit(Approx)

NOTE:Total fixed costs and variable cost per unit do not change with change in units